OpenStudy (anonymous):

okay i just need someone to tell me if my answer is correct: When the equation 3x - 4y = 24 is written in slope-intercept form, what is the value of m? i got 3/4?

OpenStudy (anonymous):

yes you have the correct answer

OpenStudy (anonymous):

can you explain to me explain to me how you got it so i can see if i did it the same way?

OpenStudy (anonymous):

@Hero is my answer correct i have a feeling she was just saying so

OpenStudy (anonymous):

or he

hero (hero):

3x - 4y = 24 3x - 24 = 4y 3/4x - 6 = y m = 3/4

OpenStudy (anonymous):

so 'm correct?

OpenStudy (anonymous):

i mean i'm

OpenStudy (texaschic101):

3x - 4y = 24 -4y = -3x + 24 y = 3/4x - 6 3/4 is your slope

OpenStudy (anonymous):

yay i got it correct thank you so much to both of you!!

OpenStudy (texaschic101):

sometimes , if your not sure about a problem, confirmation is good to get.

OpenStudy (anonymous):

ok thanl you can you help me with some other questions? @texaschic101

OpenStudy (texaschic101):

possibly...what you got

OpenStudy (anonymous):

i will give you the answers for some of them and you can correct them but i have like slopes union and things similar

OpenStudy (texaschic101):

I will do what I can

OpenStudy (anonymous):

ok thank you so much

OpenStudy (anonymous):

hope this makes sense if it doesn't i will send you a picture Jackie conducted a survey of her fellow students in the freshman class to determine what flavor of ice cream they prefer to be served at their end of the year party: vanilla, chocolate, or strawberry. She made the Venn diagram below to show her class sponsor. A Venn diagram titled Ice Cream Flavors, indicating the number of freshman students who prefer each of three flavors. Each of the three circles in the Venn diagram represents an ice cream flavor and are labeled Vanilla which is Set V, Chocolate which is Set C, and Strawberry which is Set S. The number of students to only Set V is 165, the number of students to only Set C is 278, and the number of students to only Set S is 113. The number of students for all three flavors is 37. The number of students to both Set V and Set C is 32. The number of students to both Set C and Set S is 71. The number of students to both Set S and Set V is 54. How many students belong to the set ∼C ∪ (S ∩ ∼V)?

OpenStudy (anonymous):

and i got 322

OpenStudy (anonymous):

 here if this helps this is how the graph looks like

OpenStudy (anonymous):

(S ∩ ∼V) = 54 students. 278 + 54 = 332? i might be mistaken but this is my work to back up my

OpenStudy (texaschic101):

I am sorry...I am not sure about this. You might better ask someone else....sorry

OpenStudy (anonymous):

np here's the next question

OpenStudy (anonymous):

An individual's phone number contains seven digits, not including the area code, from the set A shown below. A = {0, 1, 2, 3, 4, 5, 6, 7, 8, 9} Set B represents the digits in Brent's phone number. B = {5, 5, 5, 3, 0, 9, 9} Set C represents the digits in Charlie's phone number. C = {8, 6, 7, 5, 3, 0, 9} How many even numbers are in the set ∼(B ∩ C)?

OpenStudy (texaschic101):

what did you get ?

OpenStudy (anonymous):

2 because it's in both c and b

OpenStudy (texaschic101):

I am better with slopes then unions and intersections

OpenStudy (texaschic101):

tag hero.....he is good at everything

OpenStudy (anonymous):

ok @Hero

OpenStudy (anonymous):

@dss can you help me?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank you so much my answer 332 do you think i'm correct

OpenStudy (anonymous):

You're looking for the intersection of s and not v

OpenStudy (anonymous):

is that right

OpenStudy (anonymous):

oh i'm confused now

OpenStudy (anonymous):

OK hang on

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

∼C ∪ (S ∩ ∼V) 332 U 71=403 This is my take on the question. I don't know if I'm right.

hero (hero):

403 is correct for ∼C ∪ (S ∩ ∼V)

OpenStudy (anonymous):

Thanks Hero...I haven't done that type of equation for a long time.

OpenStudy (anonymous):

oh thank you so much